# Hypothesis Generation with Large Language Models

### v1

@@ Abstract
Effective generation of novel hypotheses is instrumental to scientific progress. So far, researchers have been the main powerhouse behind hypothesis generation by painstaking data analysis and thinking (also known as the Eureka moment). In this paper, we examine the potential of large language models (LLMs) to generate hypotheses. We focus on hypothesis generation based on data (i.e., labeled examples). To enable LLMs to handle arbitrarily long contexts, we generate initial hypotheses from a small number of examples and then update them iteratively to improve the quality of hypotheses. Inspired by multi-armed bandits, we design a reward function to inform the exploitation-exploration tradeoff in the update process. Our algorithm is able to generate hypotheses that enable much better predictive performance than few-shot prompting in classification tasks, improving accuracy by 31.7% on a synthetic dataset and by 13.9%, 3.3% and, 24.9% on three real-world datasets. We also outperform supervised learning by 12.8% and 11.2% on two challenging real-world datasets. Furthermore, we find that the generated hypotheses not only corroborate human-verified theories but also uncover new insights for the tasks.
@@ 1 Introduction
Hypothesis generation drives scientific progress. Mendel’s hypothesis on allele pairs lays the foundation for modern genetics; Einstein’s hypothesis in general theory of relativity led to the prediction and subsequent confirmation of gravitational waves. In the context of language modeling, the hypothesis on scaling law inspires recent progress in large language models (LLMs) (Chowdhery et al., 2022). Despite the importance of hypothesis generation, as Ludwig & Mullainathan (2024) points out, science has been curiously asymmetric. While many scientific publications present extensive formal and empirical evaluation of hypotheses, the generation of hypotheses happens off-stage by researchers. In order to generate novel hypotheses, researchers may read literature, analyze data, pick the brain of each other, and even “hallucinate” (see Kekulé’s discovery of the structure of the benzene molecule (Rothenberg, 1995)). Given the rise of large language models (Brown et al., 2020; Anthropic, 2023; OpenAI, 2023b), we examine their potential of providing much needed assistance in hypothesis generation in this work.
@@
To generate high-quality hypotheses with LLMs, we propose an algorithm inspired by the upper confidence bound algorithm in multi-armed bandits (Auer, 2002) (HypoGeniC, Hypothesis Generation in Context; see Figure 1). Given initial hypotheses generated from a small number of examples, we need to assess their quality and propose new hypotheses to address their deficiencies. To navigate this exploration-exploitation tradeoff, we introduce a reward function and evaluate the top $k$ hypotheses for each training example. We maintain a wrong example bank to capture the gap in knowledge of the current hypotheses pool, and generate new hypotheses based on the wrong example bank to close the gap.
@@
Figure 1: Illustration of HypoGeniC. During update stage, we evaluate the top $k$ hypotheses on each new training example and update the reward based on the prediction corretness. If the number of hypotheses that got the example wrong exceeds a certain threshold, we add the example to a wrong example bank. The wrong example bank is then used to generate new hypotheses.
@@
The generated hypotheses naturally enable an interpretable hypothesis-based classifier. We propose a suite of inference strategies given a set of hypotheses. We apply our method to one synthetic task where there is a single known valid hypothesis and three real-world tasks (Deceptive reviews, and Tweet popularity). The real-world tasks focus on deception detection and message popularity prediction, which are known to be challenging even for humans (Ott et al., 2011; Salganik et al., 2006). Our algorithm can recover the hypothesis in the synthetic task and also provide useful hypotheses for the real-world tasks. In fact, our generated hypotheses consistently outperform few-shot in-context learning baselines across all four tasks (31.7% in Shoe sales, 13.9% in Deceptive reviews, 3.3% in and 24.9% in Tweet popularity). The predictive performance matches and even outperforms oracle supervised learning with RoBERTa except in Deceptive reviews.
@@
It is important to emphasize that although the utility of hypotheses in assisting downstream classification serves as an indicator for LLMs’ ability to generate hypotheses, our primary interest lies in the quality of the hypotheses. Thus, it is critical for the hypotheses to be interpretable beyond the LLM used to produce the hypotheses. We show that hypotheses generated by one LLM (e.g., GPT-3.5-turbo) can be used to make accurate inference by another LLM (e.g., Mixtral). On an out-of-distribution dataset for Deceptive reviews, we can even outperform the oracle fine-tuned RoBERTa. Such cross generalization provides strong evidence that we are able to generate hypotheses of high quality. Furthermore, through a qualitative analysis, our generated hypotheses not only confirm theories from existing literature but also provide new insights about the task. For instance, one novel hypothesis is that “reviews that mention personal experiences or special occasions, such as birthdays, anniversaries, or weddings, are more likely to be truthful”. We encourage future research on deception detection to explore these novel hypotheses.
Our work is connected to many recent studies on using LLMs to propose “hypotheses”, notably, Qiu et al. (2024) and Zhong et al. (2023). Qiu et al. (2024) is motivated by testing the ability of LLMs to perform human-like induction reasoning, and Zhong et al. (2023) aims to support open-ended exploration. While similar in spirit, we examine the case of generating theories between input and labels for challenging problems where even researchers may struggle with proposing new hypotheses.
To summarize, we highlight the following contributions:
@@
  • We propose a novel computational framework for generating and evaluating hypotheses with large language models.
@@
  • We find our generated hypotheses to corroborate existing findings while also providing new insights for the tasks.
@@ 2 Method
We begin with a description of the problem formulation. Given a set $\mathcal{S}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ where $x_{i}$ is an example and $y_{i}$ is the corresponding label, the goal is to learn a set of hypotheses =$\mathcal{H}=\{h_{1},...,h_{m}\}$ that describe theories of relationships between $x$ and $y$. To this end, we prompt an LLM to summarize demonstration examples into high-level hypotheses (§ 2.1). Then, during inference, the LLM makes inference based on the generated hypothesis (§ 2.2). Our code is available at https://github.com/ChicagoHAI/hypothesis_generation.
@@ 2.1 Hypothesis Generation
In the update stage, for a training example $s$, we select the top $k$ high-reward hypotheses from the hypothesis bank $\mathcal{H}$. The LLM is prompted to make a prediction with each of the top $k$ high-reward hypotheses on $s$. Then we compute the accuracy of the inference and accordingly update the reward for each of the hypotheses. If $w_{hyp}$ hypotheses predict incorrectly for the example $s$, then $s$ is added to a wrong example pool $\mathcal{W}$. Once the wrong example pool reaches a max size of $w_{max}$, the wrong examples in $\mathcal{W}$ are used to generate new hypotheses, as shown in Algorithm 1. The wrong example pool represents the gap in knowledge that the current pool of hypotheses has for the dataset. Thus by generating new hypotheses, the algorithm fills in these gaps. We update $\mathcal{H}$ with the newly generated hypotheses according to the rewards.
@@
$r_{i}=\frac{\sum_{(x_{j},y_{j})\in\mathcal{S}_{i}}I(y_{i}=\hat{y}_{i})}{\mathcal{S}_{i}}+\alpha\sqrt{\frac{\log{t}}{|\mathcal{S}_{i}|}},$ (1)
@@
$\mathcal{H}$leftarrow generate_hypotheses($\{\mathcal{S}_{i}:i\leq\mathtt{num\_init}\}$) $\triangleright$ Initialize hypothesis bank
3:$\mathcal{W}$leftarrow {}
4:for $(x_{t},y_{t})\in\mathcal{S}$ :
5:   $\mathcal{H}_{\mathrm{top}}=\leftarrow\{h:\text{$h\in\mathcal{H}$ is amongst top $k$ based on reward}\}$
6:   for $h\in\mathcal{H}_{\mathrm{top}}$ :
@@
8:      update_reward($h,y_{t},\hat{y}_{t}^{h}$)
9:   if |$|\{\mathbf{wrong}(\hat{y}_{t}^{h}):h\in\mathcal{H}\}|\geq w_{hyp}$ : $\triangleright$ $w_{hyp}$ is dynamically determined, see § B.1
10:      $\mathcal{W}$leftarrow $\mathcal{W}$$\cup\{(x_{t},y_{t})\}$
11:   if $|\mathcal{W}|=w_{max}$ :
12:      $\mathcal{N}\leftarrow$ generate_hypotheses($\mathcal{W}$)
13:      $\mathcal{W}$leftarrow {}
14:      $\mathcal{H}\leftarrow\{h:\text{$h$ is amongst top $H$ based on reward in $\mathcal{H}\cup\mathcal{N}$}\}$
15:return $\mathcal{H}$
@@ 2.2 Hypothesis-based Inference
  • Best-accuracy hypothesis. The hypothesis $h$ with the highest accuracy from the hypothesis bank $\mathcal{H}$ is included in the prompt to guide the model to perform inference.
  • Filter and weighted vote. One hypothesis may not be enough to explain the data. Thus, this approach uses a combination of relevant hypotheses to make predictions for a single example. For each example, we first filter hypotheses by prompting an LLM to judge which hypotheses are relevant to the example. Next, an LLM is prompted to generate predictions for each of the relevant hypotheses, and these predictions are aggregated with weighted vote, where the weight is the training accuracy of the corresponding hypothesis.
@@ 3 Experiment Setup
Next, we introduce the experiment setup to evaluate HypoGeniC.
@@ 3.1 Tasks and Datasets
The choice of appropriate tasks is critical for evaluating the ability of LLMs to generate hypothesis. The focus of our work is on generating hypotheses based on observed data. A prerequisite is that potential hypotheses do exist. In the context of classification, it implies that the classification performance is non-trivial. In addition, we need to ensure that the hypotheses describing the data are likely not a priori known by LLMs, which rules out standard tasks such as sentiment analysis. Therefore, we use four datasets that satisfy these requirements: a synthetic task where we know the true hypothesis and three real-world datasets that exhibit complex underlying patterns and constitute widely studied social science problems.
@@
Predicting popularity is a notoriously challenging task in social sciences because it is known to be affected by seemingly random factors (Salganik et al., 2006). We use two datasets in this work: nd Tweet popularity. s derived from a dataset in the Upworthy Research Archive (Matias et al., 2021). The original dataset was collected through A/B testing, where each user was shown pairs of a headline and image for multiple packages (articles). Each user was exposed to only one of these pairs per package, and the clicks were recorded for each pair per package. [^1] The Upworthy Research Archive only provides the image IDs instead of the graphics. We thus only use the headlines for our dataset. This process resulted in a total of 150,816 headlines across 22,666 packages. We construct a binary classification dataset by choosing for each package the headline that received the most clicks and the one that received the fewest. We remove all sets of duplicate headlines, which results in our version of the ataset. The task for this dataset is to deduce which headline had more clicks in a pair. Tweet popularity uses a dataset of 13,174 tweet pairs (Tan et al., 2014), which are matched by the topic and the author. Similar to the task is to predict which one received more retweets.
@@ 3.2 Baselines, Oracles, and Evaluation Metrics
    Supervised Learning. We fine-tune RoBERTa on each of the datasets to serve as a non-interpretable oracle. We include results for training on 200 examples and 1000 examples. Since fine-tuning update model weights, we expect RoBERTa to set the upper bound on in-distribution datasets.
@@ 4 Results
To demonstrate the effectiveness of our hypothesis generation approach, we present results via three evaluation methods. First, we show that in the standard supervised learning setup, our generated hypotheses enable more accurate predictions than baselines and even oracles when using a small set of examples. Second, we evaluate the generated hypotheses by checking whether they can generalize across different inference LLMs and to out-of-distribution datasets. We find surprisingly consistent performance even when using a different LLM to make inference from the generated hypotheses. Finally, we conduct a qualitative analysis to show that the generated hypotheses not only corroborate with existing theories but also provide novel insights about the tasks at hand.
@@ 4.1 Performance on Heldout Test Sets
As discussed in the introduction, a side product of our approach is an interpretable hypothesis-based classifier. We compare its performance with standard supervised learning with fine-tuned RoBERTa and few-shot prompt learning (Table 1).
Our generated hypotheses improve inference over standard zero-shot and few-shot inference. Across all LLMs, HypoGeniC outperforms the zero-shot learning by an average of 60% on Shoe sales, 22.7% on Deceptive reviews, 5.1% on and 30.6% on Tweet popularity. Similarly, we find that HypoGeniC shows an increase from few-shot learning by 31.7% on Shoe sales, 13.9% on Deceptive reviews, 3.3% on and 24.9% on Tweet popularity. Note that these results are inflated on Tweet popularity as safety mode is triggered for Mixtral and Claude-2.1 for zero-shot and few-shot learning respectively. The results show that hypothesis-based inference can increase the performance of LLMs significantly. One exception is that our method performs slightly worse (by 1%) than the few-shot baseline in the Tweet popularity with GPT-3.5-turbo. One possible reason is that the few-shot demonstrations are effective at eliciting the pretraining knowledge in GPT-3.5-turbo, possibly due to a large amount of tweets in pretraining data. For more detailed results, refer to Appendix C.
@@
HypoGeniC matches or even exceeds RoBERTa with the same number of training examples on most datasets. Both HypoGeniC and RoBERTa yield 100% on the syntheic dataset. Moreover, HypoGeniC is 12.8% and 11.2% better than RoBERTa on nd Tweet popularity respectively with 200 training examples. Since RoBERTa learns by updating model weights to minimize the cross-entropy loss, it tends to benefit from more training examples, so we increase training examples to 1000 for RoBERTa. Despite the accuracy boost from more training examples, we find that HypoGeniC’s best result still outperforms RoBERTa by 3.7% on nd 0.7% on Tweet popularity. One exception, however, is the Deceptive reviews dataset. We suspect that as word-level features are very useful in this dataset (Ott et al., 2011), they could be tougher for LLMs to extract but easier for fine-tuned models to grasp.
@@
Updating hypothesis bank leads to hypotheses of higher quality. Comparing HypoGeniC with the “no updates” results, we find that updating hypotheses generally leads to better hypotheses, suggesting that our algorithm is effective at improving hypothesis quality. The improvement is on average 0.7% on Shoe sales, 5.8% on Deceptive reviews, 8.1% on and 7% on Tweet popularity. Another advantage of HypoGeniC over “no updates” is that sometimes the training examples exceed the context window size of LLMs, which can lead to degraded performance (Figures 4 and 3).
@@
Effect of inference strategy. Figure 2 shows HypoGeniC results with different inference strategies on Deceptive reviews. Single-step adaptive inference is the most effective. Generally, we find that hypotheses to be one-sided, focusing on either characteristics of truthful or deceptive reviews. We thus need to consider more than one hypothesis to make a correct prediction, so best-accuracy hypothesis or two-step adaptive inference would not be ideal for this dataset. On the other datasets, we find that the effect of inference strategy is much smaller. Best-accuracy hypothesis is sufficient for Shoe sales and and filter and weighted vote works best for Tweet popularity. Results for all datasets are in § C.1. Whichever inference strategy we use, the trend of HypoGeniC against few-shot learning and RoBERTa remains largely the same.
@@
Generally, having more training examples and a larger hypothesis pool improves performance. We show performance for different methods as number of training examples increase in Figures 4, 3 and 5. We find HypoGeniC accuracy steadily increases as training size increases on Shoe sales, suggesting that an LLM is more likely to generate the best hypothesis given more examples. For the real-world datasets, however, the performance sometimes peaks at training size at 25 or 100 before reaching to 200. We suspect that the evaluation of the hypothesis bank would be less stable for the real-world datasets, since more than one correct hypotheses are needed for the task. We also find that using a hypothesis pool of size 20 leads to better performance than using a pool of size 3.

### v2

@@ Abstract
Effective generation of novel hypotheses is instrumental to scientific progress. So far, researchers have been the main powerhouse behind hypothesis generation by painstaking data analysis and thinking (also known as the Eureka moment). In this paper, we examine the potential of large language models (LLMs) to generate hypotheses. We focus on hypothesis generation based on data (i.e., labeled examples). To enable LLMs to handle long contexts, we generate initial hypotheses from a small number of examples and then update them iteratively to improve the quality of hypotheses. Inspired by multi-armed bandits, we design a reward function to inform the exploitation-exploration tradeoff in the update process. Our algorithm is able to generate hypotheses that enable much better predictive performance than few-shot prompting in classification tasks, improving accuracy by 31.7% on a synthetic dataset and by 13.9%, 3.3% and, 24.9% on three real-world datasets. We also outperform supervised learning by 12.1% and 11.6% on two challenging real-world datasets. Furthermore, we find that the generated hypotheses not only corroborate human-verified theories but also uncover new insights for the tasks.
@@ 1 Introduction
Hypothesis generation drives scientific progress. Mendel’s hypothesis on allele pairs lays the foundation for modern genetics; Einstein’s hypothesis in general theory of relativity led to the prediction and subsequent confirmation of gravitational waves. In the context of language modeling, the hypothesis on scaling law inspires recent progress in large language models (LLMs) (Kaplan et al., 2020). Despite the importance of hypothesis generation, as Ludwig & Mullainathan (2024) point out, science has been curiously asymmetric. While many scientific publications present extensive formal and empirical evaluation of hypotheses, the generation of hypotheses happens off-stage by researchers. In order to generate novel hypotheses, researchers may read literature, analyze data, pick the brain of each other, and even “hallucinate” (see Kekulé’s discovery of the structure of the benzene molecule (Rothenberg, 1995)). Given the rise of large language models (Brown et al., 2020; Anthropic, 2023; OpenAI, 2023b), we examine their potential of providing much needed assistance in hypothesis generation in this work.
@@
To generate high-quality hypotheses with LLMs, we propose an algorithm inspired by the upper confidence bound algorithm in multi-armed bandits (Auer, 2002) (HypoGeniC, Hypothesis Generation in Context; see Figure 1). Given initial hypotheses generated from a small number of examples, we need to assess their quality and propose new hypotheses to address their deficiencies. To navigate this exploration-exploitation tradeoff, we introduce a reward function and evaluate the top $k$ hypotheses for each training example. We maintain a wrong example bank to capture the gap in knowledge of the hypotheses pool, and generate new hypotheses based on the wrong example bank to close the gap.
@@
Figure 1: Illustration of HypoGeniC. During update stage, we evaluate the top $k$ hypotheses on each new training example and update the reward based on the prediction correctness. If the number of hypotheses that got the example wrong exceeds a certain threshold, we add the example to a wrong example bank. The wrong example bank is then used to generate new hypotheses.
@@
The generated hypotheses naturally enable an interpretable hypothesis-based classifier. We propose a suite of inference strategies given a set of hypotheses. We apply our method to one synthetic task where there is a single known valid hypothesis and three real-world tasks (Deceptive reviews, and Tweet popularity). The real-world tasks focus on deception detection and message popularity prediction, which are known to be challenging even for humans (Ott et al., 2011; Salganik et al., 2006). Our algorithm can recover the hypothesis in the synthetic task and also provide useful hypotheses for the real-world tasks. In fact, our generated hypotheses consistently outperform few-shot in-context learning baselines across all four tasks (31.7% in Shoe sales, 13.9% in Deceptive reviews, 3.3% in and 24.9% in Tweet popularity). The predictive performance matches and even outperforms oracle supervised learning with RoBERTa and Llama-2-7B except in Deceptive reviews.
@@
It is important to emphasize that although the utility of hypotheses in assisting downstream classification serves as an indicator for LLMs’ ability to generate hypotheses, our goal is not to maximize the classification performance. Rather, our primary interest lies in the quality of the hypotheses. Thus, it is critical for the hypotheses to be interpretable beyond the LLM used to produce the hypotheses. We show that hypotheses generated by one LLM (e.g., GPT-3.5-turbo) can be used to make accurate inference by another LLM (e.g., Mixtral). On an out-of-distribution dataset for Deceptive reviews, we can even outperform the oracle fine-tuned RoBERTa. Such cross generalization provides strong evidence that we are able to generate hypotheses of high quality. Furthermore, through a qualitative analysis, our generated hypotheses not only confirm theories from existing literature but also provide new insights about the tasks. For instance, one novel hypothesis is that “reviews that mention personal experiences or special occasions, such as birthdays, anniversaries, or weddings, are more likely to be truthful”. We encourage future research on deception detection to explore these novel hypotheses.
To summarize, we make the following contributions:
@@
  • We propose a novel computational framework for generating and evaluating hypotheses with LLMs.
@@
  • Our generated hypotheses corroborate existing findings while also providing new insights for the tasks.
We have publicly released the code and data for HypoGeniC at https://github.com/ChicagoHAI/hypothesis-generation.
@@ 2 Method
We begin with a description of the problem formulation. Given a set $\mathcal{S}=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ where $x_{i}$ is an example and $y_{i}$ is the corresponding label, the goal is to learn a set of hypotheses =$\mathcal{H}=\{h_{1},...,h_{m}\}$ that describe theories of relationships between $x$ and $y$. To this end, we prompt an LLM to summarize demonstration examples into high-level hypotheses (§ 2.1). Then, during inference, the LLM makes inference based on the generated hypothesis (§ 2.2).
@@ 2.1 Hypothesis Generation
In the update stage, for a training example $s$, we select the top $k$ high-reward hypotheses from the hypothesis bank $\mathcal{H}$. The LLM is prompted to make a prediction with each of the top $k$ high-reward hypotheses on $s$. Then we compute the accuracy of the inference and accordingly update the reward for each of the hypotheses. If $w_{hyp}$ hypotheses predict incorrectly for the example $s$, then $s$ is added to a wrong example pool $\mathcal{W}$. Once the wrong example pool reaches a max size of $w_{max}$, the wrong examples in $\mathcal{W}$ are used to generate new hypotheses. The wrong example pool represents the gap in knowledge that the current pool of hypotheses has for the dataset. Thus, by generating new hypotheses, the algorithm fills in these gaps. We update $\mathcal{H}$ with the newly generated hypotheses as per the rewards.
@@
$r_{i}=\frac{\sum_{(x_{j},y_{j})\in\mathcal{S}_{i}}I(y_{j}=\hat{y}_{j})}{|\mathcal{S}_{i}|}+\alpha\sqrt{\frac{\log{t}}{|\mathcal{S}_{i}|}},$ (1)
@@
2:// Initialize hypothesis bank
$\mathcal{H}$leftarrow generate_hypotheses($\{\mathcal{S}_{i}:i\leq\mathtt{num\_init}\}$)
4:$\mathcal{W}$leftarrow {}
5:for $(x_{t},y_{t})\in\mathcal{S}$ :
6:   $\mathcal{H}_{\mathrm{top}}\leftarrow\{h:\text{$h\in\mathcal{H}$ has top k reward}\}$
7:   for $h\in\mathcal{H}_{\mathrm{top}}$ :
@@
9:      update_reward($h,y_{t},\hat{y}_{t}^{h}$)
10:   if |$|\{\mathbf{wrong}(\hat{y}_{t}^{h}):h\in\mathcal{H}\}|\geq w_{hyp}$ :
11:      // $w_{hyp}$ is dynamically determined, see § B.1
12:      $\mathcal{W}$leftarrow $\mathcal{W}$$\cup\{(x_{t},y_{t})\}$
13:   if $|\mathcal{W}|=w_{max}$ :
14:      $\mathcal{N}\leftarrow$ generate_hypotheses($\mathcal{W}$)
15:      $\mathcal{W}$leftarrow {}
16:      $\mathcal{H}\leftarrow\{h:\text{$h\in\mathcal{H}\cup\mathcal{N}$ has top k reward}\}$
17:return $\mathcal{H}$
@@ 2.2 Hypothesis-based Inference
  • Best-accuracy hypothesis. The hypothesis $h$ with the highest accuracy from the hypothesis bank is included in the prompt to guide the model to perform inference.
  • Filter and weighted vote. One hypothesis may not be enough to explain the data. Thus, this approach uses a combination of relevant hypotheses to make predictions for a single example. We first filter hypotheses by prompting an LLM to judge which hypotheses are relevant to the example. Next, an LLM is prompted to generate predictions for each of the relevant hypotheses, and these predictions are aggregated with weighted vote, where the weight is the training accuracy of the corresponding hypothesis.
@@ 3 Experiment Setup
We introduce the experiment setup to evaluate HypoGeniC.
@@ 3.1 Tasks and Datasets
The choice of appropriate tasks is critical for evaluating the ability of LLMs to generate hypothesis. The focus of our work is on generating hypotheses based on observed data. A prerequisite is that potential hypotheses do exist. In the context of classification, it implies that the classification performance is non-trivial. In addition, we need to ensure that the hypotheses describing the data are likely not a priori known by LLMs, which rules out standard tasks such as sentiment analysis. Therefore, we use four datasets that satisfy these requirements: a synthetic task with a known true hypothesis and three real-world datasets that exhibit complex underlying patterns and constitute widely studied social science problems.
@@
Predicting popularity is a notoriously challenging task in social sciences because it is known to be affected by seemingly random factors (Salganik et al., 2006). We use two datasets in this work: nd Tweet popularity. s derived from a dataset in the Upworthy Research Archive (Matias et al., 2021). The original dataset was collected through A/B testing, where each user was shown pairs of a headline and image for multiple packages (articles). Each user was exposed to only one of these pairs per package, and the clicks were recorded for each pair per package. [^1] The Upworthy Research Archive only provides the image IDs instead of the graphics. We thus only use the headlines for our dataset. This process resulted in a total of 150,816 headlines across 22,666 packages. We construct a binary classification dataset by choosing the headlines that received the most clicks and least clicks for each package. We remove all sets of duplicate headlines, which results in our version of the ataset. The task for this dataset is to deduce which headline had more clicks in a pair. Tweet popularity uses a dataset of 13,174 tweet pairs (Tan et al., 2014), which are matched by the topic and the author. Similar to the task is to predict which one received more retweets.
@@ 3.2 Baselines, Oracles, and Evaluation Metrics
    Supervised Learning. We fine-tune RoBERTa (Liu et al., 2019) and Llama-2-7B (Touvron et al., 2023) on each of the datasets to serve as a non-interpretable oracle. We include results for training on 200 examples and 1000 examples. Since fine-tuning update model weights, we expect RoBERTa and Llama-2-7B to set the upper bound on in-distribution datasets.
@@ 4 Results
To demonstrate the effectiveness of our hypothesis generation approach, we present results via three evaluation methods. First, we show that in the standard supervised learning setup, our generated hypotheses enable more accurate predictions than baselines and even oracles when using a small set of examples. Second, we evaluate the generated hypotheses by checking whether they can generalize across different inference LLMs and to out-of-distribution datasets. We find surprisingly consistent performance even when using a different LLM to make inference from the generated hypotheses. So, we conduct a qualitative analysis to show that the generated hypotheses not only corroborate existing theories but also provide novel insights about the tasks at hand.
@@ 4.1 Performance on Heldout Test Sets
As discussed in the introduction, a side product of our approach is an interpretable hypothesis-based classifier. We compare its performance with standard supervised learning with the fine-tuned models and few-shot in-context learning (Table 1).
Our generated hypotheses improve inference over standard zero-shot and few-shot inference.
Across all LLMs, HypoGeniC outperforms the zero-shot learning by an average of 60% on Shoe sales, 22.7% on Deceptive reviews, 5.1% on and 30.6% on Tweet popularity. Similarly, we find that HypoGeniC shows an increase from few-shot learning by 31.7% on Shoe sales, 13.9% on Deceptive reviews, 3.3% on and 24.9% on Tweet popularity. Note that these results are inflated on Tweet popularity as safety mode is triggered for Mixtral and Claude-2.1 for zero-shot and few-shot learning respectively. After computing the 95% confidence intervals (with a binomial distribution assumption) for our results, the following results are significant for the real life datasets: HypoGeniC for Deceptive reviews and Tweet popularity with Claude-2.1 and Mixtral, when comparing to their respective few shot baselines. If we relax the confidence interval to 90%, the result for ith Mixtral is also statistically significant. These results demonstrate that hypothesis-based inference can increase the performance of LLMs significantly. Further results can be found in Table 5. One exception is that our method performs slightly worse (by 1%) than the few-shot baseline in the Tweet popularity with GPT-3.5-turbo. One possible reason is that the few-shot demonstrations are effective at eliciting the pretraining knowledge in GPT-3.5-turbo, possibly due to a large amount of tweets in pretraining data. More detailed results are in Appendix C.
@@
HypoGeniC matches or even exceeds the fine-tuned models with the same number of training examples on most datasets.
@@
Both HypoGeniC and the fine-tuned models yield 100% on the syntheic dataset. Moreover, HypoGeniC is 12.8% and 11.2% better than RoBERTa, and 12.1% and 11.6% better than Llama-2-7B, on nd Tweet popularity respectively with 200 training examples. Since the fine-tuned models learns by updating model weights to minimize the cross-entropy loss, it tends to benefit from more training examples, so we increase training examples to 1000 for the fine-tuned models. Despite the accuracy boost from more training examples, we find that HypoGeniC’s best result still outperforms RoBERTa by 3.7% and 0.7%, and Llama-2-7B by 3.7% and 11.4%, on nd Tweet popularity, respectively. One exception, however, is the Deceptive reviews dataset. We suspect that as word-level features are very useful in this dataset (Ott et al., 2011), they could be tougher for LLMs to extract but easier for fine-tuned models to grasp.
Updating the hypothesis bank leads to hypotheses of higher quality.
@@
Comparing HypoGeniC with the “no updates” results, we find that updating hypotheses generally leads to better hypotheses, suggesting that our algorithm is effective at improving hypothesis quality. The improvement is on average 0.7% on Shoe sales, 5.8% on Deceptive reviews, 8.1% on and 7% on Tweet popularity. Another advantage of HypoGeniC over “no updates” is that sometimes the training examples exceed the context window size of LLMs, which can lead to degraded performance (Figures 4 and 3).
Effect of inference strategy.
@@
Figure 2 shows HypoGeniC results with different inference strategies on Deceptive reviews. Single-step adaptive inference is the most effective. Generally, we find hypotheses to be one-sided, focusing on either characteristics of truthful or deceptive reviews. We thus need to consider more than one hypothesis to make a correct prediction, so best-accuracy hypothesis or two-step adaptive inference are not ideal. On the other datasets, we find that the effect of inference strategy is much smaller (§ C.1). Best-accuracy hypothesis is sufficient for Shoe sales and and filter and weighted vote works best for Tweet popularity. Whichever inference strategy we use, the trend of HypoGeniC against few-shot learning and the fine-tuned models remains largely the same.
Generally, having more training examples and a larger hypothesis pool improves performance.
@@
We show performance for different methods as number of training examples increase in Figures 4, 3 and 5. We find HypoGeniC accuracy steadily increases as training size increases on Shoe sales, suggesting that an LLM is more likely to generate the best hypothesis given more examples. For the real-world datasets, however, the performance sometimes peaks at training size at 25 or 100 before reaching to 200. We suspect that the evaluation of the hypothesis bank would be less stable for the real-world datasets, since more than one correct hypotheses are needed for the task. We also find that using a hypothesis pool of size 20 leads to better performance than using a pool of size 3.
@@
Although this classification experiment is convenient to run and demonstrates that our generated hypotheses are reasonable, our main goal is to generate high-quality hypotheses rather than maximizing the performance of this particular way of using the hypotheses. The next two experiments are essential in understanding the quality of hypotheses through generalization and manual analysis.